# Imitate the Good and Avoid the Bad: An Incremental Approach to Safe Reinforcement Learning

### 16 Dec 2023 v1

@@ 1 Introduction
It’s important to note that imitation within our context differs from the conventional IL approaches from the aforementioned works. Here, our approach not only involves mimicking the behavior of “good” demonstrations but also actively avoiding the bad ones. To the best of our knowledge, this marks the first time the concept of learning to avoid bad demonstrations is introduced within the realm of IL. Additionally, in a standard IL algorithm, the set of expert demonstrations is fixed beforehand. In contrast, in our context, the set of demonstrations is generated by a pre-trained or learning policy, thus allowing it to expand as training progresses. These factors collectively pave the way for the development of a novel IL algorithm that is well-suited to our specific context.
@@ 2.2 Imitation Learning
BC has a strong theoretical foundation, as it can guarantee to return the exact expert policy $\pi^{E}$, i.e., $\pi^{E}$ is the unique optimal solution to BC. However, BC ignores environmental dynamics and only works with offline learning, so it often requires a huge number of samples to achieve a desired performance (Ross, Gordon, and Bagnell 2011).
@@ 3.2 Theoretical Insights
For any $\lambda>0$, if there exists a policy $\pi^{*}$ such that $P_{\pi^{*}}(\tau)=0$ for all $\tau\in\Omega^{B}$, and
then $\pi^{*}$ is an optimal policy to (1).
@@ 4 Self-Imitation based Safe RL
     # Update the “good” and “bad” sets, $\sigma$ is the deviation
     $R_{B}\leftarrow\mathbb{E}_{\tau\sim T}\left[R(\tau)\right]-\sigma_{\tau\sim T}[R(\tau)]$
@@ 5 EXPERIMENTS
We conduct experiments to compare our method against some state-of-the-art Constrained RL algorithms: FOCOPS (Zhang, Vuong, and Ross 2020), CUP (Yang et al. 2022), CPO (Achiam et al. 2017) [^3] FOCOPS, CUP, and CPO implementations can be found on https://github.com/PKU-Alignment/omnisafe.. For the sake of completeness, we also include PPO-Lagrangian  (Ray, Achiam, and Amodei 2019) and unconstrained PPO. We use PPO-Lagrangian to train our pre-trained policy and name our algorithm as SIM, standing for Self-IMitation based safe RL algorithm.
Through the experiments, we aim to address the following questions.
  • (Q1)
    Would SIM outperform state-of-the-art constrained RL algorithms?
  • (Q2)
    Is it necessary to use both good and bad demonstrations in the training?
  • (Q3)
    How is SIM compared to a BC-based and GAIL-based algorithm?
  • (Q4)
    How does SIM perform with different expertise levels of the initial policy $\pi^{0}$? Can it benefit from a not well-trained policy
In the following, we provide experiments to address Q1-Q4. Additional experiments addressing Q5-Q8 can be found in the appendix. We set the cost limit as $c_{max}=18$. We, however, train the initial policy $\pi^{0}$ with a higher cost limit of $c_{max}=28$, which allows us to generate more trajectories of high rewards and more unsafe trajectories. We test our method on 6 SafetyGym environments (Ji et al. 2023). We also simplify the names of environments, e.g., SafetyPointGoal1-v0 is renamed as SafetyPointGoal.
@@ Appendix C Additional Experiments
In this section, we provide experiments to answer 4 additional questions:
@@ C.1 Hyper-parameter selection
Moreover, in practice, instead of solving the theoretical problem (4), our implementation applies a more stable version with the contribution of the Chi-square divergence $\chi^{2}$:
@@ C.5 Conditional Value at Risk
C.6 Low-quality Initial Policy
@@
C.7 Experiments on Mujoco domains
@@
Eer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 13 Mar 2024 v3

@@ 1 Introduction
It is important to note that imitation within our context differs from the conventional IL approaches from the aforementioned works. Here, our approach not only involves mimicking the behavior of “good” demonstrations but also actively avoiding the bad ones. To the best of our knowledge, this marks the first time the concept of learning to avoid bad demonstrations is introduced within the realm of IL. Additionally, in a standard IL algorithm, the set of expert demonstrations is fixed beforehand. In contrast, in our context, the set of demonstrations is generated by a pre-trained or learning policy, thus allowing it to expand as training progresses. These factors collectively pave the way for the development of a novel IL algorithm that is well-suited to our specific context.
Contrastive Learning: Our framework is also related to the context of Contrastive Learning (CL). CL was first introduced by (Bromley et al. 1993) with the Siamese architecture to create a mapping function for the inputs into a target space where two similar samples should be close while two different classes should be far away. There are several famous applications of CL in computer vision (Noroozi and Favaro 2016; He et al. 2019; Grill et al. 2020), natural language processing (Clark et al. 2020; Gao, Yao, and Chen 2021), recommendation systems (Zhou et al. 2021; Xie et al. 2021), and reinforcement learning (Fu et al. 2021; Laskin, Srinivas, and Abbeel 2020). Our algorithm shares a similar spirit with CL and also marks the first time the idea of contrastive learning being applied in IL.
@@ 2.2 Imitation Learning
BC has a strong theoretical foundation but ignores environmental dynamics and only works with offline learning, requiring a huge number of samples to achieve a desired performance (Ross, Gordon, and Bagnell 2011).
@@ 3.2 Theoretical Insights
For any $\lambda>0$, if there exists a policy $\pi^{*}$ such that $P_{\pi^{*}}(\tau)=0$ for all $\tau\in\Omega^{B}$, and $P_{\pi^{*}}(\tau)=\frac{P_{\pi^{0}}(\tau)}{\sum_{\tau^{\prime}\in\Omega^{G}}P_{\pi^{0}}(\tau^{\prime})};~{}\forall\tau\in\Omega^{G}$ then $\pi^{*}$ is an optimal policy to (1).
@@ 4 Self-Imitation based Safe RL
     # Update the “good” and “bad” sets
     $R_{B}\leftarrow\mathbb{E}_{\tau\sim T}\left[R(\tau)\right]-\sigma_{\tau\sim T}[R(\tau)]$, #$\sigma$ is the deviation
@@ 5 EXPERIMENTS
We conduct experiments to compare our method against some state-of-the-art Constrained RL algorithms: FOCOPS (Zhang, Vuong, and Ross 2020), CUP (Yang et al. 2022), CPO (Achiam et al. 2017) [^3] FOCOPS, CUP, and CPO implementations can be found on https://github.com/PKU-Alignment/omnisafe.. For the sake of completeness, we also include PPO-Lagrangian  (Ray, Achiam, and Amodei 2019) and unconstrained PPO. We use PPO-Lagrangian to train our pre-trained policy and name our algorithm as SIM, standing for Self-IMitation based safe RL algorithm. Through the following experiments, we aim to address the following questions: (Q1) Would SIM outperform state-of-the-art constrained RL algorithms? (Q2) Is it necessary to use both good and bad demonstrations in the training? (Q3) How is SIM compared to a BC-based and GAIL-based algorithm? (Q4) How does SIM perform with different expertise levels of the initial policy $\pi^{0}$? Can it benefit from a not well-trained policy?
We set the cost limit as $c_{max}=18$. We, however, train the initial policy $\pi^{0}$ with a higher cost limit of $c_{max}=28$, which allows us to generate more trajectories of high rewards and more unsafe trajectories. We test our method on 6 SafetyGym environments (Ji et al. 2023). We also simplify the names of environments, e.g., SafetyPointGoal1-v0 is renamed as SafetyPointGoal.
@@ 6 CONCLUSION
Acknowledgment
This research/project is supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-RP-2020-016).
@@ Appendix C Additional Experiments
In this section, we provide experiments to answer 5 additional questions:
@@
  • (Q9)
    Do the number of initial good trajectories impact to the final performance?
@@ C.1 Hyper-parameter selection
Moreover, to enhance stability, we use a Chi-square function $\phi(x)=x-\frac{1}{a}x^{2}$ to regularize the loss function in (4) :
@@ C.5 Conditional Value at Risk
C.6 Number of initial expert demonstrations
In this section, we aim to address the impact of the number of initial expert trajectories on our final performance Q9. To this end, we run the experiments with different numbers of expert trajectories, taken from the set $[5,10,25,50,100,300,500]$. The detailed results are shown in Figure 16. Here, it is easy to observe that having a large number of expert demonstrations can offer better performance. This is possibly because having this high number of expert demonstrations can reduce the number of explorations in the environment and quickly understand the criteria for classifying good and bad trajectories.
SafetyPointGoal
SafetyCarGoal
Return
Cost
SafetyPointGoal
SafetyCarGoal
Figure 16: Best performance of the 7 different number of expert trajectories.
C.7 Low-quality Initial Policy
@@
C.8 Experiments on Mujoco domains